You move a charge q from ri = 9.1 m to rf = 4.6 m from a fixed charge Q = 10.7 nC. What is the difference in potential for these two positions?




1
Expert's answer
2021-12-03T12:44:09-0500

given properties:

"r_1=9.1m"

"r_2=4.6m"

"q=10.7nC = 10.7\u00b710^{-9}C"

"k = 9·10^9\frac{N·m^2}{C^2}"

"\Delta\varphi - ?"

difference in potential for these two positions is "\Delta\varphi=k·\large\frac{q}{r_1-r_2}" "= \large\frac{9·10^9·10.7·10^{-9}}{4.5}" "=21.4V"